OpenStudy (anonymous):

What is the volume of a sphere with a radius of 6 inches? 112Ï€ cubic inches 288Ï€ cubic inches 290Ï€ cubic inches 300Ï€ cubic inches

OpenStudy (igreen):

\(\sf V = \dfrac{4}{3} \pi r^3\)

OpenStudy (igreen):

Just ignore pi and plug in the radius. \(\sf V = \dfrac{4}{3}(6^3)\) Simplify

OpenStudy (swamijatin2):

don't solve the the pi 4/3x216=? solve it

OpenStudy (anonymous):

is that 4 over 3

OpenStudy (anonymous):

C

OpenStudy (anonymous):

hello at @iGreen could u help